# Perceptually Optimized Super Resolution

### v1

@@ 4.2 Perceptual Contrast Modeling
$C_{t}(f,p)=\frac{\text{sign}\left(C_{n}(f,p)\right)\cdot\left|C_{n}(f,p)\right|^{\alpha}}{1+\frac{1}{|N|}\sum_{q\in N(p)}\left|C_{n}(f,q)\right|^{\beta}}$ (6)
Here, the numerator models the effect of self-masking, while the denominator models the effect of spatial masking from the $5\times 5$ neighborhood in the same band. If the neighboring pixels/region have contrast on similar spatial frequencies, the effective perceptual contrast is reduced due to masking. $\alpha,\beta\in[0,1]$ are parameters of our model controlling masking modeling. The value used in our model are $\alpha=0.7$ and $\beta=0.2$.
@@
4.3 Perceptual Optimization
Given an input image or an image patch, the goal of the perceptual optimization is to find the maximum attenuation that is tolerable by the HVS. We model the attenuation curve using Gaussian fall-off:
$\alpha^{\prime}(f)=\frac{1}{a\sqrt{2\pi}}\exp\left(-\frac{(f-b)^{2}}{2a^{2}}\right)+c$ (7)
where $f$ is the spatial frequency and $a,b,c$ are parameters that we wish to estimate. According to the contrast model, the attenuation remains undetectable by an observer if the contrast difference between the original image and the attenuated one is under 1 JND. Consequently, to find the attenuation that results in maximum performance gains yet imperceptible quality loss, we optimize the attenuation curve such that it results in exactly 1 JND difference, i.e.,
$\forall_{f}\,\,C_{t}(p,f)-C_{t}^{\prime}(p,f)=1,$ (8)
where $C_{t}(p,f)$ represents the perceptual contrast of the input image patch and $C_{t}^{\prime}(p,f)$ represents the perceptual contrast of the network output. Assuming the the attenuation curves are a response of the network/branch to the input, the attenuation is a modulation of the physical image contrast at different frequencies.
$\alpha^{\prime}(f)=\frac{C^{\prime}(f,p)}{C(f.p)}=\frac{C^{\prime}_{n}(f,p)}{C_{n}(f.p)}$ (9)
where $C_{n}(f,p)$ is the CSF multiplied input contrast encoded by our contrast pyramid, and our immediate goal is to estimate the tolerable output contrast $C^{\prime}_{n}(f,p)$ as constrained by Eq. 7. We aim to estimate values $a$, $b$, and $c$ for which the following equality holds:
$\frac{\text{sign}\left(C_{n}(f,p)\right)\cdot\left|C_{n}(f,p)\right|^{\alpha}}{1+\frac{1}{|N|}\displaystyle\sum_{q\in N(p)}\left|C_{n}(f,q)\right|^{\beta}}-\frac{\text{sign}\left(C_{n}^{\prime}(f,p)\right)\cdot\left|C_{n}^{\prime}(f,p)\right|^{\alpha}}{1+\frac{1}{|N|}\displaystyle\sum_{q\in N(p)}\left|C_{n}^{\prime}(f,q)\right|^{\beta}}=1$ (10)
It is evident that $C_{n}^{\prime}(f,q)$ cannot be directly calculated from this equation due to the visual spatial masking term present in the denominator. Consequently, we make the assumption that the contrast masking for the up-sampled output patch can be approximated by that of the input patch. Furthermore, assuming that the sign of the contrast remains unaltered throughout the neural network processing Eq. 10 can be simplified to:
$\frac{\text{sign}\left(C_{n}(f,p)\right)\cdot\left(\left|C_{n}(f,p)\right|^{\alpha}-\left|C_{n}^{\prime}(f,p)\right|^{\alpha}\right)}{1+\frac{1}{|N|}\displaystyle\sum_{q\in N(p)}\left|C_{n}(f,q)\right|^{\beta}}=1$ (11)
From this equation, we can derive $C_{n}^{\prime}(f,p)$ directly as follows:
$C_{n}^{\prime}(f,p)=\left|\left|C_{n}^{\prime}(f,p)\right|^{\alpha}-\left(1+\sum_{q\in N(p)}\frac{\left|C_{n}^{\prime}(f,q)\right|^{\beta}}{|N|}\right)\right|^{1/\alpha}$ (12)
$t_{1}=\frac{C^{\prime}(f_{1},p)}{C(f_{1},p)},t_{2}=\frac{C^{\prime}(f_{2},p)}{C(f_{2},p)},t_{3}=\frac{C^{\prime}(f_{3},p)}{C(f_{3},p)}$ (13)
The required values of $a$, $b$, and $c$ can be estimated by fitting the attenuation model from Eq. 7 to the three data points: $\alpha^{\prime}(f_{i})=t_{i}$ for $i\in\{1,2,3\}$. Subsequently, the network/branch that will produce the most closely approximating attenuation for the given patch is selected. The selection of the branch is made in accordance with the following procedure:
$\text{branch}=\max_{i\in[0,...,n]}\left\{\frac{t\cdot\hat{t}_{i}}{\lVert t\rVert\lVert\hat{t}_{i}\rVert}\right\}$ (14)
where $t$ is the vector of the computed attenuations and $\hat{t}_{i}$ is the vector of the estimated attenuation produced by Eq. 7.
4.4 Model Efficiency
In our application, it is essential that the overhead of the perceptual model is minimal compared to the performance improvements provided by the SR method. Although we prototyped our solution using Python and evaluated it on a single CPU thread, it has been shown that a similar processing pipeline can be implemented efficiently on a GPU. This is primarily due to the nature of the computation, which consists of independent per-pixel operations (see Eq. 12). Additionally, the construction of the contrast pyramid can be accelerated using MIPMAP functionality. Rencently, [37] have demonstrated a perceptual model on the Meta Quest 2 VR headset, achieving 2K resolution with a single execution time of under 1 ms.
5 Experimental Setup
In this section, we go through the methodology to validate the efficacy of our approach.
5.1 Validation SR Models
@@
As mentioned in Section. 4, we employ our perceptual model to optimize SR for two cases. The first approach adds earlier exit points (branches) to the original network. Using earlier points, i.e., shallower branches, leads to less computation and lower reconstruction quality. To test this case, we employ the popular and seminal VDSR [18] network. In the case of VDSR, a neural network with 19 branched outputs was created as per the setting in Fig. 3 (top). After each ReLU activation function, an exit point was added, structured identically to the final layer of the original network., similar to [16]. Our task was to to use the perceptual model to select the appropriate branch (per image patch) such that there is no noticeable quality loss.
@@
The second approach is reducing the depth or number of channels of a network to make it more efficient. In this case, we have a number of candidate networks with different number of channels-per-layer, and our task is to employ our perceptual model to select the appropriate one per image patch. To test this case, we employ the EDSR [23] network. In the case of EDSR, five networks with varying numbers of channels per layer (256, 128, 64, 16, 8) were trained independently, the training procedure was the same as described in [23]. Our task was to use the perceptual model to select the appropriate network (per image patch) such that there is no noticeable quality loss.
We selected the two candidate networks because of their popularity and robustness and demonstrated efficacy over the years. Our algorithm is network agnostic as all SR models will follow similar attenuation characteristics, which are grounded in the nature of natural images and the Fourier nature of the SR problem. So to the best of out knowledge, we do not foresee any issues for its application to other SR models.
To evaluate the effectiveness of our method, we tested the model on images and videos, as our model is additionally capable of processing information about the temporal frequencies present in videos.
5.2 Input Patch Size
The input patch size (size of patches into which the image is divided, as shown in Fig. 3) to our perceptual model is equivalent to the size of the receptive field of the upsampling model employed. In the case of the VDSR network, the input patch is $\frac{40}{k}\times\frac{40}{k}$ pixels. This is due to the fact that, during the perceptual model prediction, the low-resolution image is being considered. Before being conveyed to the VDSR network, the LR image is upsampled through bicubic interpolation. Consequently, in the event of $\times$4 upsampling, the input to the perceptual model is 10x10 pixels, corresponding to a patch of 40x40 pixels in the image upsampled with bicubic interpolation. In the case of the EDSR network, the input patch is 48 $\times$ 48 pixels, taken from the low-resolution image. As no prior upsampling is involved in the input image, there is no need to consider a lower-size patch. In certain instances, utilizing input patches smaller than the neural network’s receptive field may be advantageous, particularly in the context of smaller images.
5.3 Subjective Quality Study
Setup For our experiments, we assumed standard office conditions where the content is viewed on a 27-inch Dell U2723QE display with a resolution of 3840 $\times$ 2160 and a peak luminance of 400 $cd/m^{2}$ from the viewing distance of 60 cm. The perceptual model was initialized according to this setup. During our experiments with human subjects, the viewing distance was controlled with the use of a chin rest that allowed to maintain constant viewing conditions throughout the experiments for all participants.
@@ 5.3 Subjective Quality Study
Task We employed a 2AFC experimental protocol. The task was a forced choice between two test cases in relation to a given high quality reference. Users were instructed to select the test case which was perceptually more similar to the reference. On the right half of the screen was always the high quality reference, and the left had either case A or B in randomized order. Users could use the space-bar to switch between A and B, and press ENTER when they had made their choice. The two test cases were
@@
  • A: the output of our perceptually optimized algorithm that selects the appropriate network/branch for the corresponding image patch.
@@
  • B: The output of the unaltered full deep network applied to the whole image.
@@
Stimuli We employed 24 high quality natural scenes for our user study. The scenes were selected for diversity in characteristics such as luminance, contrast and texture.
@@ 6.1 Quantitative Results
The quantitative results for the image datasets are presented in Tab. 1. The proposed method allows for comparable performance to the original networks while reducing the computational cost. For instance, the $\times$2 and $\times$4 upsampling operations exhibit a reduced computational cost ranging from 58% to 22% and from 70% to 20%, respectively.
As anticipated, greater savings achieved with $\times$2 upsampling in comparison to $\times$4 upsampling is in line with expectations. This is due to the increased presence of high-frequency information in the images, which allows for reconstructing certain parts of the image with less computational power.
@@
Bicubic ×$\times 2$ 32.32 0.923 28.60 0.859 28.22 0.834 25.48 0.840 31.23 0.898
VDSR ×$\times 2$ 34.15 0.946 89.60G (100%) 29.98 0.899 173.68G (100%) 29.68 0.885 131.04G (100%) 27.17 0.900 144.99G (100%) 32.65 0.931 2.04T (100%)
VDSR-ours ×$\times 2$ 32.53 0.936 51.99G (58%) 29.29 0.895 100.18G (57%) 29.40 0.883 61.63G (47%) 26.48 0.889 85.43G (58%) 32.08 0.928 1.06T (51%)
EDSR ×$\times 2$ 36.42 0.954 1.54T (100%) 32.03 0.905 2.97T (100%) 30.62 0.888 2.25T (100%) 30.42 0.932 2.56T (100%) 34.84 0.939 31.64T (100%)
EDSR-ours ×$\times 2$ 36.00 0.951 353.09G (22%) 31.58 0.901 691.09G (23%) 30.26 0.882 466.39G (20%) 29.74 0.925 606.62G (23%) 34.25 0.925 7.27T (22%)
Bicubic ×$\times 4$ 26.98 0.790 24.28 0.676 24.54 0.638 21.89 0.642 26.80 0.756
VDSR ×$\times 4$ 28.13 0.827 89.60G (100%) 25.01 0.709 173.68G (100%) 25.11 0.670 131.04G (100%) 22.75 0.698 572.38G (100%) 27.52 0.786 2.04T (100%)
VDSR-ours ×$\times 4$ 27.77 0.823 64.13G (71%) 24.86 0.713 118.64G (68%) 25.08 0.670 69.27G (52%) 22.65 0.699 387.01G (67%) 27.44 0.788 1.29T (63%)
EDSR ×$\times 4$ 30.60 0.878 579.49G (100%) 26.95 0.753 1.00T (100%) 26.01 0.706 695.39G (100%) 24.82 0.776 3.16T (100%) 29.05 0.822 10.32T (100%)
EDSR-ours ×$\times 4$ 30.27 0.874 145.34G (25%) 26.69 0.750 236.71G (23%) 25.82 0.701 173.54G (24%) 24.48 0.765 746.13G (23%) 28.75 0.816 2.46T (23%)
@@ 6.2 Subjective Quality Results
Metrics such as PSNR, SSIM etc are not representative of the intricacies of how humans perceive image quality. As our perceptual model is based on robust and detailed modeling of early human vision, out hypothesis such that the is little noticeable quality loss can only be aptly verified using subjective quality studies.

### v2

@@ 4.3 Optimization
$\frac{\text{sign}\left(C^{\prime}_{n}(f,p)\right)\cdot\left|C^{\prime}_{n}(f,p)\right|^{\alpha}}{1+\frac{1}{|N|}\displaystyle\sum_{q\in N(p)}\left|C^{\prime}_{n}(f,q)\right|^{\beta}}-\frac{\text{sign}\left(C_{n}(f,p)\right)\cdot\left|C_{n}(f,p)\right|^{\alpha}}{1+\frac{1}{|N|}\displaystyle\sum_{q\in N(p)}\left|C_{n}(f,q)\right|^{\beta}}=1$,
where the numerators encode the CSF-weighted contrast values, while the denominators model visual masking effect. It is evident that $C_{n}^{\prime}(f,q)$ cannot be directly calculated from this equation due to the visual masking term present in the denominator. To address this problem, similarly to the work of [46], we make the assumption that the contrast masking for the up-sampled output patch can be approximated by that of the input patch. Furthermore, knowing that the sign of the contrast should remain unaltered throughout the neural network processing Eq. 6 we can derive $C_{n}^{\prime}(f,p)$ directly as:
$C_{n}^{\prime}(f,p)=\left|\left(1+\sum_{q\in N(p)}\frac{\left|C_{n}(f,q)\right|^{\beta}}{|N|}\right)-\left|C_{n}(f,p)\right|^{\alpha}\right|^{1/\alpha}$
$t_{i}=\frac{C_{n}^{\prime}(f_{i},p)}{C_{n}(f_{i},p)},i\in{1,2,3}$ (8)
It is important to note that thanks to the derivation in Eq. 7, the attenuation can be computed directly from $C_{n}$, i.e., the input patch only.
Having the tolerable attenuation $t=\{t_{1},t_{2},t_{3}\}$, we can find the most suitable SR network/branch by identifying the one with most similar attenuation curve (Section 4.1, Eq. 3). More formally, this step can be defined using following optimization problem:
$\text{branch/network}=\arg\max_{j}\left\{\frac{t\cdot\hat{t}_{j}}{\lVert t\rVert\lVert\hat{t}_{j}\rVert}\right\},$ (9)
where $\hat{t}_{j}=\{\alpha^{\prime}(f_{1}),\alpha^{\prime}(f_{2}),\alpha^{\prime}(f_{3})\}$ is the vector of the estimated attenuation produced by Eq. 3 for a given netwrok/brach $j$.
4.4 Model Efficiency
In our application, it is essential that the overhead of the HVPF is minimal compared to the performance improvements provided by the SR method. Although we prototyped our solution using Python and evaluated it on a single CPU thread, it has been shown that a similar processing pipeline can be implemented efficiently on a GPU. This is primarily due to the nature of the computation, which consists of independent per-pixel operations (see Eq. 7). Additionally, the construction of the contrast pyramid can be accelerated using MIPMAP functionality. Rencently,  [45] have demonstrated a HVPF on the Meta Quest 2 VR headset, achieving 2K resolution with a single execution time of under 1 ms.
5 Experimental Setup
@@
Validation on SR Models
@@
As mentioned in Section. 4, we employ our HVPF to optimize SR for two cases. The first approach adds earlier exit points (branches) to the original network. Using earlier points, i.e., shallower branches, leads to less computation and lower reconstruction quality. To test this case, we employ the popular and seminal VDSR [22] network. A neural network with 19 branched outputs was created as per the setting in Fig. 3 (top). After each ReLU activation function, an exit point was added, structured identically to the final layer of the original network., similar to [20]. Our task was to to use the HVPF to select the appropriate branch (per image patch) such that there is no noticeable quality loss.
@@
The second approach is reducing the depth or number of channels of a network to make it more efficient. In this case, we have a number of candidate networks with different number of channels-per-layer, and our task is to employ our HVPF to select the appropriate one per image patch. To test this case, we employ the EDSR [28] network. In the case of EDSR, five networks with varying numbers of channels per layer (256, 128, 64, 16, 8) were trained independently, the training procedure was the same as described in [28]. Our task was to use the HVPF to select the appropriate network (per image patch) such that there is no noticeable quality loss. We include additional details on our choice of patch sizes in Sec. D.
On the generality of our visual framework.
We selected EDSR and VDSR models as test benchmarks for HVPF due to their widespread use, robustness and demonstrated efficacy over the years. Our algorithm is network agnostic: most SR approaches will follow similar attenuation characteristics, which are grounded in the nature of natural images and the Fourier nature of the SR problem [33]. In Fig. 4 we show attenuation curves for Transformer-based models, where increasing model size directly correlates with its ability to reconstruct higher spatial frequency content. This is also in line with previous research on implicit models for vision and graphics, where increasing the number of weights directly correlates with the capacity of the model to learn higher frequency content [34] and together with our results on CNNs makes us confident on the generality of our HVPF to be leveraged with any SR method.
Figure 4: Attenuation curves derived from the SwinIR, a transformer based SR model. As the size of the model increases the capability of reconstructing higher frequency details increases.
Subjective Quality Study.
We developed a perceptual experiment with human subjects in order to validate our approach.
@@
Task. We employed a 2AFC experimental protocol. The task was a forced choice between two test cases in relation to a given high quality reference. Users were instructed to select the test case which was perceived more similar to the reference. On the right side we display the high quality reference; while the left side displayed either
@@
  • A) our HVPF-powered SR, selecting the appropriate network/branch for the corresponding image patch or,
@@
  • B) The output of the unaltered full deep network applied to the whole image.
These were displayed in randomized order. Users could use the space-bar to switch between A and B, and press ENTER when they had made their choice.
@@
Stimuli. We employed 24 high-quality natural scenes for our user study. The scenes were selected for diversity in characteristics such as luminance, contrast, and texture. Each scene was downscaled by a factor of $\times 4$ and upscaled back to the original resolution.
In Sec. A we included additional details on the experimental setup.
@@ 6 Results and Discussion
To evaluate the effectiveness of our method, we tested the model on images and videos, as our model is additionally capable of handling the temporal frequencies present in videos.
@@ 6.1 Quantitative Results
The quantitative results for the image datasets are presented in Tab. 1. The proposed method allows for comparable performance to the original networks while reducing the computational cost. For instance, the $\times$2 and $\times$4 upsampling operations exhibit a reduced computational cost ranging from 58% to 22% and from 70% to 20%, respectively. For a single patch of size $10\times 10$ the computational cost of our method is 39KFlops, while for a patch size of $35\times 35$ is 477KFlops.
Greater savings were achieved with $\times$2 upsampling in comparison to $\times$4 upsampling, which is in line with expectations. This is due to the increased presence of high-frequency information in the images, which allows for reconstructing certain parts of the image with less computational power.
@@
In Sec. C we provide an explanation for the choice of FLOPS as a measure of efficiency.
@@
Bicubic $\times 2$ 32.32 0.923 28.60 0.859 28.22 0.834 25.48 0.840 31.23 0.898
VDSR $\times 2$ 34.15 0.946 89.60G (100%) 29.98 0.899 173.68G (100%) 29.68 0.885 131.04G (100%) 27.17 0.900 144.99G (100%) 32.65 0.931 2.04T (100%)
VDSR-ours $\times 2$ 32.53 0.936 51.99G (58%) 29.29 0.895 100.18G (57%) 29.40 0.883 61.63G (47%) 26.48 0.889 85.43G (58%) 32.08 0.928 1.06T (51%)
EDSR $\times 2$ 36.42 0.954 1.54T (100%) 32.03 0.905 2.97T (100%) 30.62 0.888 2.25T (100%) 30.42 0.932 2.56T (100%) 34.84 0.939 31.64T (100%)
EDSR-ours $\times 2$ 36.00 0.951 353.09G (22%) 31.58 0.901 691.09G (23%) 30.26 0.882 466.39G (20%) 29.74 0.925 606.62G (23%) 34.25 0.925 7.27T (22%)
Bicubic $\times 4$ 26.98 0.790 24.28 0.676 24.54 0.638 21.89 0.642 26.80 0.756
VDSR $\times 4$ 28.13 0.827 89.60G (100%) 25.01 0.709 173.68G (100%) 25.11 0.670 131.04G (100%) 22.75 0.698 572.38G (100%) 27.52 0.786 2.04T (100%)
VDSR-ours $\times 4$ 27.77 0.823 64.13G (71%) 24.86 0.713 118.64G (68%) 25.08 0.670 69.27G (52%) 22.65 0.699 387.01G (67%) 27.44 0.788 1.29T (63%)
EDSR $\times 4$ 30.60 0.878 579.49G (100%) 26.95 0.753 1.00T (100%) 26.01 0.706 695.39G (100%) 24.82 0.776 3.16T (100%) 29.05 0.822 10.32T (100%)
EDSR-ours $\times 4$ 30.27 0.874 145.34G (25%) 26.69 0.750 236.71G (23%) 25.82 0.701 173.54G (24%) 24.48 0.765 746.13G (23%) 28.75 0.816 2.46T (23%)
@@ 6.2 Subjective Quality Results
It is well-known that metrics such as PSNR, SSIM, etc are not correlated with human quality assessments, unable to model the intricacies of how humans perceive image quality [16, 38, 35]. As our HVPF is based on a robust and detailed modeling of early visual processing in a scene dependent manner (unlike heuristic metrics trained over a large set of images e.g [58]), our hypothesis (little noticeable quality loss) can only be aptly verified using subjective quality studies.
@@
A more detailed study on the network channel depth SR alternative (EDSR, bottom Fig. 3) is presented in Sec. B.1, providing similar conclusions. In summary, the studies demonstrate that the results indeed conform with our hypothesis that there is no perceivable loss in visual quality, even when there are significant computational savings through the application of our HVPF.
Finally, we also include promising early results on foveated SR (super resolution aware of gaze position, leveraging the substantially degraded visual acuity on the peripheral vision) in Fig. 6 and in Sec. E.
Figure 6: Our model predictions based on gaze position with $\times 4$ super-resolution. In the first column, we have the original image and the corresponding quality map. In the other column we have the quality maps at different gaze positions.
7 Limitations and Future work
Although the method demonstrated satisfactory performance with regard to video content, it was observed that aliasing issues were present. This phenomenon is common when individually processing frames as isolated still images. In addition, our method does not explicitly guarantee the spatial consistency between the patches. However, no such inconsistencies were noted or reported by the participants in our experiments. It would be of interest to investigate the potential of our HVPF in conjunction with video-specific SR techniques. Evaluating the performance of our model within foveation-aware frameworks is also an interesting avenue for future work, given that our choice of contrast-sensitivity function (StelaCSF) is capable of modelling contrast sensitivity based on eccentricity. Finally, our estimation is based solely on image luminance; we could further leverage additional information, such as colour, by integrating recent advances in color-aware CSFs like CastleCSF [3].
8 Conclusions
We present a thorough novel framework to leverage the specific deficiencies of the HVS to optimize computational resources in the context of SR. HVPF is fast, robust, and can be seamlessly integrated into any SR framework, adaptively optimizing computational resources to the areas in the image that require it, from the assumption that a human will be the final observer of the resulting SR image. While normally SR methods are evaluated in terms of reconstruction quality, through metrics such as PSNR or SSIM, these metrics do not model the intricacies of human visual quality assessment, and have been repeatedly demonstrated as uncorrelated with human visual perception [16, 38, 35]. In contrast, we validate our framework through a series of human studies, showcasing indistinguishable quality at substantially reduced computational costs. Our results demonstrate that even with FLOP reductions by factors of-and-greater than 3x, our HVPF minimally degrades the SR output such that the final result is not visually distinguishable from the output of a fully-resourced deep network. We are confident that our method can be further extended in the future to better integrate with video-specific SR approaches, and will be particularly relevant on VR and AR eccentricity-aware SR frameworks, where computational savings can be pushed significantly further.
Acknowledgement
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grand agreement No 804226).
Supplementary Material
A Subjective Quality Study Setup
Setup. For our experiments, we assumed standard office conditions where the content is viewed on a 27-inch Dell U2723QE display with a resolution of 3840 $\times$ 2160 and a peak luminance of 400 $cd/m^{2}$, from a viewing distance of 60 cm. All our results are calculated according to this setup. During our experiments with human subjects, the viewing distance was controlled with the use of a chin rest that allowed to maintain constant viewing conditions throughout the experiments for all participants.
B Additional Subjective Quality Results
B.1 Channel Depth Application
@@
B.2 Video Content
Setup and Task The same experimental protocol as described in Sec. 5 was employed for this experiment.
Stimuli The stimuli were derived from seven natural videos from the Inter4k [44] dataset.
Data preparation Each video frame was down-sampled by a factor of eight and then up-sampled by a factor of four using the VDSR network, specifically the network branching application.